# Fully non-linear elliptic equations on complex manifolds

### v2

@@ 1. Introduction
where $\lambda[\omega^{-1}(\chi+\sqrt{-1}\partial\overline{\partial}u)]=(\lambda_{1},\cdots,\lambda_{n})$ denotes the $n$-tuple of eigenvalues of $\chi+\sqrt{-1}\partial\overline{\partial}u$ with respect to $\omega$. Henceforth, $f$ is a smooth symmetric function defined in an open symmetric convex cone $\Gamma\subset\mathbb{R}^{n}$ with vertex at origin and $\partial\Gamma\neq\emptyset$,
@@ 1.1. Equations on closed manifolds
The main issue is to establish first and second order estimates. In order to set up first order estimate, we use the blow-up method that was developed by Dinew-Kołodziej [16] in the case of complex $k$-Hessian equations. The argument is based on a Liouville type theorem and a second order estimate of the form
@@
Such a second estimate was first established by Hou-Ma-Wu [37] for complex $k$-Hessian equations on closed Kähler manifolds. More recently, their works were extended by Székelyhidi [52] to general fully nonlinear elliptic equations satisfying (1.6). For more related work, we refer to [55, 56, 53] for Monge-Ampère equation for $(n-1)$-PSH functions and also to [68] for complex $k$-Hessian equation.
@@
In addition, $\mathring{\Gamma}_{\mathcal{G}}^{f}$ denotes the interior of $\Gamma_{\mathcal{G}}^{f}$. And then we may prove that $v$ is a $\mathring{\Gamma}_{\mathcal{G}}^{f}$-solution in the sense of [52, Definition 15]. This will contradict to Székelyhidi’s Liouville type theorem [52, Theorem 20] (in analogy with Dinew-Kołodziej [16, Theorem 3.2]) and then give the desired gradient bound.
Comparing with the literature mentioned above, in the proof of Theorem 1.3 a new difficulty we face is that the positive lower bound $\sum_{i=1}^{n}f_{i}\geqslant\delta$ possibly fails when dropping (1.6). In order to overcome this, we propose Lemma 3.3 below. The idea of exploiting such type lemma is from Guan [25]. With the aid of this lemma, we can make use of $(\underline{u}-u)$ in the construction of auxiliary function in order to overcome the difficulties caused by geometric quantities and equations. As usual, the casewise argument of Hou-Ma-Wu [37] is employed. This trick is motivated by the treatment of Chou-Wang [12] on real Hessian equation. In addition, we follow the technique of Szekélyhidi [52] to treat the troublesome third order terms arising from nontrivial torsion tensors of non-Kähler metrics.
@@
Remark 1.5.
As shown by Lemma 3.5, the unbound and concavity of $f$ together imply (1.3). Thus we do not impose the monotonicity condition in Theorem 1.4.
@@ 1.2. The Dirichlet problem
Below, we establish the quantitative boundary estimate for general equations.
@@
(1.14) $\inf_{M}\psi=\sup_{\partial\Gamma}f,$
(1.15) $f\in C^{\infty}(\Gamma)\cap C(\bar{\Gamma}),\mbox{ where }\bar{\Gamma}=\Gamma\cup\partial\Gamma,$
which will be shown in Sections 9 and 10.
@@ 3.1. A lemma for subsolution
Then there is a constant $\varepsilon$ depending on $\sigma$, $\mu$, $\delta$, $R$ such that if $\lambda\in\partial\Gamma^{\sigma}$, then either
(3.2) $\sum_{i=1}^{n}f_{i}(\lambda)(\mu_{i}-\lambda_{i})\geqslant\varepsilon\left(1+\sum_{i=1}^{n}f_{i}(\lambda)\right)$
@@
(3.3) $f_{i}(\lambda)\geqslant\varepsilon\left(1+\sum_{j=1}^{n}f_{j}(\lambda)\right),\,\forall 1\leqslant i\leqslant n.$
@@
Our proof is only based on the concavity assumption. Denote $\tilde{\mu}=\mu-\delta\vec{\bf 1}$. By (3.1), there exists a positive constant $R_{0}$ depending only on $\tilde{\mu}$ and $R$ such that
@@
where $\vec{\bf e}_{i}$ is the $i$-th standard basis vector in $\mathbb{R}^{n}$. Denote $\delta_{0}=\underset{1\leqslant i\leqslant n}{\min}\{f(\tilde{\mu}+R_{0}\vec{\bf e}_{i})-\sigma\}$.
To see this, we pick some $R_{0}>0$ such that $\tilde{\mu}_{i}+R_{0}>R$, $\forall 1\leqslant i\leqslant n.$ Suppose by contradiction that $\tilde{\mu}+R_{0}\vec{\bf e}_{i^{\prime}}\in\mathbb{R}^{n}\setminus\Gamma^{\sigma}$ for some $i^{\prime}$. Then we take suitable small constant $0<\epsilon<R_{0}$ such that $\tilde{\mu}+(R_{0}-\epsilon)\vec{\bf e}_{i^{\prime}}\in\mathbb{R}^{n}\setminus\bar{\Gamma}^{\sigma}$ and $\tilde{\mu}_{i^{\prime}}+R_{0}-\epsilon>R$. From this we know that there is some $t>0$ such that $\tilde{\mu}+(R_{0}-\epsilon)\vec{\bf e}_{i^{\prime}}+t\vec{\bf 1}\in\partial\Gamma^{\sigma}$, which contradicts to (3.1).
Below we prove that the conclusion holds for $\varepsilon=\min\{\frac{\delta}{2},\frac{\delta}{2R_{0}},\frac{\delta_{0}}{2},\frac{\delta_{0}}{2R_{0}}\}.$ Without loss of generality, assume $\lambda_{1}\leqslant\cdots\leqslant\lambda_{n}$. Then $f_{1}(\lambda)\geqslant\cdots\geqslant f_{n}(\lambda).$ If
$f_{n}(\lambda)>\frac{\delta_{0}}{2R_{0}}+\frac{\delta}{2R_{0}}\sum_{i=1}^{n}f_{i}(\lambda),$
then (3.3) holds. From now on we assume $\frac{\delta_{0}}{2}+\frac{\delta}{2}\sum_{i=1}^{n}f_{i}(\lambda)\geqslant R_{0}f_{n}(\lambda).$ By the concavity of $f$, we have
$\sum_{i=1}^{n}f_{i}(\lambda)(\mu_{i}-\lambda_{i})=$ $\sum_{i=1}^{n}f_{i}(\lambda)(\tilde{\mu}_{i}-\lambda_{i})+\delta\sum_{i=1}^{n}f_{i}(\lambda)$
$\geqslant$ $f(\tilde{\mu}+R_{0}\vec{\bf e}_{n})-f(\lambda)-R_{0}f_{n}(\lambda)+\delta\sum_{i=1}^{n}f_{i}(\lambda)$
$\geqslant$ $\frac{\delta_{0}}{2}+\frac{\delta}{2}\sum_{i=1}^{n}f_{i}(\lambda).$
@@
In contrast to previous works [25, 26, 52], Lemma 3.1 stands out prominently due to its inclusion of constant terms in both (3.2) and (3.3). This feature enables us to handle general equations more simply under optimal conditions.
@@
Suppose that there exists a $\mathcal{C}$-subsolution $\underline{u}\in C^{2}(\bar{M})$. Then there exists a positive constant $\varepsilon$ such that we either have
(3.5) $F^{i\bar{j}}(\underline{\mathfrak{g}}_{i\bar{j}}-\mathfrak{g}_{i\bar{j}})\geqslant\varepsilon F^{i\bar{j}}g_{i\bar{j}}+\varepsilon$
@@
Combining Lemma 3.1 and (2.2), we can complete the proof.
@@ 3.3. Lemmas for blow-up argument
From (3.8), for any $\lambda\in\Gamma$ and $\mu\in\Gamma_{\mathcal{G}}^{f}$, $\sum_{i=1}^{n}f_{i}(\lambda)\mu_{i}\geqslant 0.$ Combining (3.7), we obtain (3.9).
@@ 3.4. A quantitative lemma
One of key ingredients for quantitative boundary estimate for double normal derivative is how to follow the track of the eigenvalues of $\left(\mathfrak{g}_{i\bar{j}}\right)$ as $\mathfrak{g}_{n\bar{n}}$ tends to infinity. To this end, we prove the following lemma. (Since the proof is somewhat complicated and long, we will leave it to Section 11).
@@ 8.1. Double normal derivative case
Applying Lemma 3.11, immediately we have
@@
When $f$ further satisfies (1.6), the unbounded condition (1.8) imposed in Theorems 1.7 and 1.8 can be replaced by a slightly broader condition
@@ 8.2. Tangential-Normal derivatives case
$w\leqslant 0\mbox{ in }\Omega_{\delta},$
$\mathcal{L}w=F^{i\bar{j}}(\mathfrak{\underline{g}}_{i\bar{j}}-\mathfrak{g}_{i\bar{j}})+2NF^{i\bar{j}}\sigma_{i}\sigma_{\bar{j}}+(2N\sigma-t)\mathcal{L}\sigma\mbox{ in }\Omega_{\delta}.$
@@
$|(u-\varphi)_{t^{\alpha}}|\leqslant C|z|\mbox{ on }\partial M\cap\bar{\Omega}_{\delta},\mbox{ }\forall 1\leqslant\alpha<2n.$
@@
Combining Cauchy-Schwarz inequality, we can prove
@@
By straightforward computations and [25, Proposition 2.19], we have
@@
In order to treat the terms involving $\sum f_{i}|\lambda_{i}|$, applying a trick used in [26] and Cauchy-Schwarz inequality, we can derive
(8.17) $\sum f_{i}|\lambda_{i}|\leqslant\epsilon\sum_{i\neq r}f_{i}\lambda_{i}^{2}+(|\underline{\lambda}|+\epsilon^{-1})\sum f_{i}+\sum f_{i}(\underline{\lambda}_{i}-\lambda_{i}),\,\forall\epsilon>0.$
@@
(8.18) $\mathcal{L}\widetilde{\Psi}\geqslant$ $(A_{1}\sqrt{K}-A_{3}C_{\Phi}-C_{1}^{\prime})F^{i\bar{j}}(\mathfrak{\underline{g}}_{i\bar{j}}-\mathfrak{g}_{i\bar{j}})+2A_{1}N\sqrt{K}F^{i\bar{j}}\sigma_{i}\sigma_{\bar{j}}$
@@
$-(C_{1}^{\prime}+A_{3}C_{\Phi})-\left(C_{1}^{\prime}+(C_{1}^{\prime}+A_{3}C_{\Phi})|\underline{\lambda}|\right)\sum_{i=1}^{n}f_{i}.$
@@
The discussion can be divided into two cases according to Lemma 3.3.
@@
$F^{i\bar{j}}(\underline{\mathfrak{g}}_{i\bar{j}}-\mathfrak{g}_{i\bar{j}})\geqslant\varepsilon F^{i\bar{j}}g_{i\bar{j}}+\varepsilon.$
@@
$F^{i\bar{j}}\sigma_{i}\sigma_{\bar{j}}\geqslant\frac{\varepsilon}{16}\left(1+\sum_{i=1}^{n}f_{i}\right)\mbox{ on }\Omega_{\delta}$
@@
$\mathcal{L}\widetilde{\Psi}\geqslant$ $\frac{A_{1}N\varepsilon}{8}\sqrt{K}\left(1+\sum_{i=1}^{n}f_{i}\right)-(C_{1}^{\prime}+A_{3}C_{\Phi})-\left(C_{1}^{\prime}+(C_{1}^{\prime}+A_{3}C_{\Phi})|\underline{\lambda}|\right)\sum_{i=1}^{n}f_{i}$
$-\left(A_{2}+A_{3}C_{\Phi}+\frac{A_{1}\varepsilon}{32}+4(C_{1}^{\prime}+A_{3}C_{\Phi})^{2}\right)\sqrt{K}\sum_{i=1}^{n}f_{i}>0$

### v3

@@ 1. Introduction
where $\lambda[\omega^{-1}(\chi+\sqrt{-1}\partial\overline{\partial}u)]=(\lambda_{1},\cdots,\lambda_{n})$ denotes the $n$-tuple of eigenvalues of $\chi+\sqrt{-1}\partial\overline{\partial}u$ with respect to $\omega$. Henceforth, $f$ is a smooth symmetric function defined in an open symmetric convex cone $\Gamma\subset\mathbb{R}^{n}$, with vertex at the origin and $\partial\Gamma\neq\emptyset$,
@@ 1.1. Equations on closed manifolds
The main issue is to establish first and second order estimates. In order to set up first order estimate, we use the blow-up method that was developed by Dinew-Kołodziej [16] in the case of complex $k$-Hessian equations. The argument is based on a Liouville type theorem and the second order estimate of the form
@@
Such a second estimate was first established by Hou-Ma-Wu [37] for complex $k$-Hessian equations on closed Kähler manifolds. More recently, their works were extended by Székelyhidi [52] to general fully nonlinear elliptic equations satisfying (1.6). For more related work, we refer to [55, 56, 53] for Monge-Ampère equation for $(n-1)$-PSH functions, and also to [68] for complex $k$-Hessian equation.
@@
In addition, $\mathring{\Gamma}_{\mathcal{G}}^{f}$ denotes the interior of $\Gamma_{\mathcal{G}}^{f}$. And then we may employ Lemma 3.7 to prove that $v$ is a $\mathring{\Gamma}_{\mathcal{G}}^{f}$-solution in the sense of [52, Definition 15]. This will contradict to Székelyhidi’s Liouville type theorem [52, Theorem 20] (in analogy with Dinew-Kołodziej [16, Theorem 3.2]) and then give the desired gradient bound.
Comparing with the literature mentioned above, in the proof of Theorem 1.3 a new difficulty we face is that the positive lower bound $\sum_{i=1}^{n}f_{i}\geqslant\delta$ possibly fails when dropping (1.6). In order to overcome this, we propose Lemmas 3.1 and 3.3 below. The idea of exploiting such type lemma is from Guan [25]. With the aid of this lemma, we can make use of $(\underline{u}-u)$ in the construction of auxiliary function in order to overcome the difficulties caused by geometric quantities and equations. The newly introduced terms of the form $1+\big{|}\sum f_{i}\lambda_{i}\big{|}$ in Lemma 3.1 will greatly simplify the process of managing and solving general equations under optimal conditions. As usual, the casewise argument of Hou-Ma-Wu [37] is employed. This trick is motivated by the treatment of Chou-Wang [12] on real Hessian equation. In addition, we follow the technique of Szekélyhidi [52] to treat the troublesome third order terms arising from nontrivial torsion tensors of non-Kähler metrics.
@@ 1.2. The Dirichlet problem
Below, based on Lemma 3.12 we establish the quantitative boundary estimate.
@@
Remark 1.8.
As shown by Lemma 3.5, the unbound and concavity of $f$ together imply (1.3). Thus we do not impose (1.3) in Theorems 1.4, 1.6 and 1.7.
@@
(1.14) $\inf_{M}\psi=\sup_{\partial\Gamma}f.$
which will be shown in Sections 9 and 10. In this case we further assume
(1.15) $f\in C^{\infty}(\Gamma)\cap C(\bar{\Gamma}),\mbox{ where }\bar{\Gamma}=\Gamma\cup\partial\Gamma.$
@@ 3.1. A lemma for subsolution
Then there is a constant $\varepsilon$ depending on $\sigma$, $\mu$, $\delta$, $R$, $f$ (explicitly see (3.7)) such that if $\lambda\in\partial\Gamma^{\sigma}$, then either
(3.2) $\sum_{i=1}^{n}f_{i}(\lambda)(\mu_{i}-\lambda_{i})\geqslant\varepsilon\left(1+\sum_{i=1}^{n}f_{i}(\lambda)+\big{|}\sum_{i=1}^{n}f_{i}(\lambda)\lambda_{i}\big{|}\right),$
@@
(3.3) $f_{i}(\lambda)\geqslant\varepsilon\left(1+\sum_{j=1}^{n}f_{j}(\lambda)+\big{|}\sum_{j=1}^{n}f_{j}(\lambda)\lambda_{j}\big{|}\right),\,\forall 1\leqslant i\leqslant n.$
@@
Our proof is based only on the concavity assumption. Denote $\tilde{\mu}=\mu-\delta\vec{\bf 1}$. By (3.1), there exists a positive constant $R_{0}$ depending only on $\tilde{\mu}$ and $R$ such that
@@
where $\vec{\bf e}_{i}$ is the $i$-th standard basis vector in $\mathbb{R}^{n}$.
To see this, we pick some $R_{0}>0$ such that $\tilde{\mu}_{i}+R_{0}>R$, $\forall 1\leqslant i\leqslant n.$ Suppose by contradiction that $\tilde{\mu}+R_{0}\vec{\bf e}_{i^{\prime}}\in\mathbb{R}^{n}\setminus\Gamma^{\sigma}$ for some $i^{\prime}$. Then we take suitable (small) constant $0<\epsilon<R_{0}$ such that $\tilde{\mu}+(R_{0}-\epsilon)\vec{\bf e}_{i^{\prime}}\in\mathbb{R}^{n}\setminus\bar{\Gamma}^{\sigma}$ and $\tilde{\mu}_{i^{\prime}}+R_{0}-\epsilon>R$. From this we know that there is some $t>0$ such that $\tilde{\mu}+(R_{0}-\epsilon)\vec{\bf e}_{i^{\prime}}+t\vec{\bf 1}\in\partial\Gamma^{\sigma}$, which contradicts to (3.1).
From (3.4) we can see that there is a small constant $0<\epsilon_{1}\ll 1$ depending only on $\tilde{\mu}$ and $\delta$ such that
(3.5) $f((1+\epsilon_{1})\tilde{\mu}+R_{0}\vec{\bf e}_{i})>\sigma,\,f((1-\epsilon_{1})\tilde{\mu}+R_{0}\vec{\bf e}_{i})>\sigma,\,\forall 1\leq i\leq n.$
Fix such $\epsilon_{1}$, we take
(3.6) $\frac{1}{1+c}=\begin{cases}1-\epsilon_{1}\,&\mbox{ if }\sum_{i=1}^{n}f_{i}(\lambda)\lambda_{i}\geq 0,\\ 1+\epsilon_{1}\,&\mbox{ otherwise.}\end{cases}$
Denote $\delta_{0}=\min\left\{\underset{1\leqslant i\leqslant n}{\min}\{f((1+\epsilon_{1})\tilde{\mu}+R_{0}\vec{\bf e}_{i})-\sigma\},\underset{1\leqslant i\leqslant n}{\min}\{f((1-\epsilon_{1})\tilde{\mu}+R_{0}\vec{\bf e}_{i})-\sigma\}\right\}.$
Below we prove that the conclusion holds for
(3.7) $\varepsilon=\min\left\{\frac{\delta_{0}}{2R_{0}},\frac{\delta(1-\epsilon_{1})}{2R_{0}},\frac{\epsilon_{1}}{2R_{0}},\frac{\delta_{0}}{2(1+\epsilon_{1})},\frac{\delta}{2},\frac{\epsilon_{1}}{2(1+\epsilon_{1})}\right\}.$
Without loss of generality, assume $\lambda_{1}\leqslant\cdots\leqslant\lambda_{n}$. Then $f_{1}(\lambda)\geqslant\cdots\geqslant f_{n}(\lambda).$ If
$f_{n}(\lambda)>\frac{\delta_{0}}{2R_{0}}+\frac{\delta}{2R_{0}(1+c)}\sum_{i=1}^{n}f_{i}(\lambda)+\frac{c}{2R_{0}(1+c)}\sum_{i=1}^{n}f_{i}(\lambda)\lambda_{i},$
then by (3.6)
$f_{n}(\lambda)>\frac{\delta_{0}}{2R_{0}}+\frac{\delta(1-\epsilon_{1})}{2R_{0}}\sum_{i=1}^{n}f_{i}(\lambda)+\frac{\epsilon_{1}}{2R_{0}}\big{|}\sum_{i=1}^{n}f_{i}(\lambda)\lambda_{i}\big{|}.$
From now on we assume
$\frac{\delta_{0}}{2}+\frac{\delta}{2(1+c)}\sum_{i=1}^{n}f_{i}(\lambda)+\frac{c}{2(1+c)}\sum_{i=1}^{n}f_{i}(\lambda)\lambda_{i}\geqslant R_{0}f_{n}(\lambda).$
Note $\mu=\tilde{\mu}+\delta\vec{\bf 1}.$ By the concavity of $f$, using (3.10) below we have
$\frac{1}{1+c}\sum_{i=1}^{n}f_{i}(\lambda)(\mu_{i}-\lambda_{i})=$ $\sum_{i=1}^{n}f_{i}(\lambda)(\frac{1}{1+c}\tilde{\mu}_{i}-\lambda_{i})+\frac{\delta}{1+c}\sum_{i=1}^{n}f_{i}(\lambda)+\frac{c}{1+c}\sum_{i=1}^{n}f_{i}(\lambda)\lambda_{i}$
$=$ $\sum_{i=1}^{n-1}f_{i}(\lambda)(\frac{1}{1+c}\tilde{\mu}_{i}-\lambda_{i})+f_{n}(\lambda)(\frac{1}{1+c}\tilde{\mu}_{n}+R_{0}-\lambda_{n})$
$-R_{0}f_{n}(\lambda)+\frac{\delta}{1+c}\sum_{i=1}^{n}f_{i}(\lambda)+\frac{c}{1+c}\sum_{i=1}^{n}f_{i}(\lambda)\lambda_{i}$
$\geqslant$ $f(\frac{1}{1+c}\tilde{\mu}+R_{0}\vec{\bf e}_{n})-f(\lambda)+\frac{\delta}{1+c}\sum_{i=1}^{n}f_{i}(\lambda)$
$+\frac{c}{1+c}\sum_{i=1}^{n}f_{i}(\lambda)\lambda_{i}-R_{0}f_{n}(\lambda)$
$\geqslant$ $\delta_{0}+\frac{\delta}{1+c}\sum_{i=1}^{n}f_{i}(\lambda)+\frac{c}{1+c}\sum_{i=1}^{n}f_{i}(\lambda)\lambda_{i}-R_{0}f_{n}(\lambda)$
$\geqslant$ $\frac{\delta_{0}}{2}+\frac{\delta}{2(1+c)}\sum_{i=1}^{n}f_{i}(\lambda)+\frac{c}{2(1+c)}\sum_{i=1}^{n}f_{i}(\lambda)\lambda_{i}.$
That is
$\sum_{i=1}^{n}f_{i}(\lambda)(\mu_{i}-\lambda_{i})\geqslant\,$ $\frac{\delta_{0}(1+c)}{2}+\frac{\delta}{2}\sum_{i=1}^{n}f_{i}(\lambda)+\frac{(1+c)\epsilon_{1}}{2}\big{|}\sum_{i=1}^{n}f_{i}(\lambda)\lambda_{i}\big{|}$
[…]
@@
In contrast to previous works [25, 26, 52], Lemma 3.1 stands out prominently due to its inclusion of the new terms of the form $1+\big{|}\sum f_{i}\lambda_{i}\big{|}$ in both (3.2) and (3.3). This feature enhances the scope of the results.
@@
Let $(\bar{M},\omega)$ be a compact Hermitian manifold of complex dimension $n\geqslant 2$ (possibly with boundary). Suppose that there exists a $\mathcal{C}$-subsolution $\underline{u}\in C^{2}(\bar{M})$ to the equation (1.1). Then there exists a positive constant $\varepsilon$, independent of $(\delta_{\psi,f})^{-1}$, such that either
(3.8) $F^{i\bar{j}}(\underline{\mathfrak{g}}_{i\bar{j}}-\mathfrak{g}_{i\bar{j}})\geqslant\varepsilon(1+F^{i\bar{j}}g_{i\bar{j}}+\big{|}F^{i\bar{j}}\mathfrak{g}_{i\bar{j}}\big{|})$
@@
Combining Lemma 3.1, (2.2) and $F^{i\bar{j}}\mathfrak{g}_{i\bar{j}}=\sum f_{i}(\lambda)\lambda_{i}$, we have either (3.8) or (3.9). From the choice of $\varepsilon$ in (3.2), we see that the constant does not depend on $(\delta_{\psi,f})^{-1}.$
@@ 3.3. Lemmas for blow-up argument
From (3.11), for any $\lambda\in\Gamma$ and $\mu\in\Gamma_{\mathcal{G}}^{f}$, we have $\sum_{i=1}^{n}f_{i}(\lambda)\mu_{i}\geqslant 0.$ Combining (3.10), we obtain (3.12).
@@ 3.4. A quantitative lemma
One of key ingredients for quantitative boundary estimate for double normal derivative is how to follow the track of the eigenvalues of $\left(\mathfrak{g}_{i\bar{j}}\right)$ as $\mathfrak{g}_{n\bar{n}}$ tends to infinity. To this end, we propose the following lemma. (Since the proof is somewhat complicated and long, we will leave it to Section 11).
@@ 8.1. Double normal derivative case
When $f$ satisfies (1.6), applying Lemma 3.11 we have immediately
@@
When $f$ further satisfies (1.6), the unbounded condition (1.8) imposed in Theorems 1.6 and 1.7 can be replaced by a slightly broader condition
@@ 8.2. Tangential-Normal derivatives case
We will highlight that the expression $\big{|}\sum f_{i}\lambda_{i}\big{|}$ in Lemma 3.1 offers a simplified method for managing the problematic terms that incorporate $\sum f_{i}|\lambda_{i}|$. This enhances some of the techniques mentioned earlier.
@@
(8.16) $w\leqslant 0\mbox{ in }\Omega_{\delta},$
(8.17) $\mathcal{L}w=F^{i\bar{j}}(\mathfrak{\underline{g}}_{i\bar{j}}-\mathfrak{g}_{i\bar{j}})+2NF^{i\bar{j}}\sigma_{i}\sigma_{\bar{j}}+(2N\sigma-t)\mathcal{L}\sigma\mbox{ in }\Omega_{\delta}.$
@@
(8.18) $|(u-\varphi)_{t^{\alpha}}|\leqslant C|z|\mbox{ on }\partial M\cap\bar{\Omega}_{\delta},\mbox{ }\forall 1\leqslant\alpha<2n.$
@@
Note that $u_{x_{n}}=2u_{n}+\sqrt{-1}u_{y_{n}}$ and $u_{x_{n}}=2u_{\bar{n}}-\sqrt{-1}u_{y_{n}}$. Combining Cauchy-Schwarz inequality, we can prove
@@
By straightforward computations and [25, Proposition 2.19], there is an index $r$ with $1\leqslant r\leqslant n$ such that
@@
In order to treat the terms involving $\sum f_{i}|\lambda_{i}|$, applying Cauchy-Schwarz inequality, we can derive
(8.20) $\sum f_{i}|\lambda_{i}|=$ $\sum_{i=1}^{n}f_{i}\lambda_{i}-2\sum_{\lambda_{i}<0}f_{i}\lambda_{i}=2\sum_{\lambda_{i}\geq 0}f_{i}\lambda_{i}-\sum_{i=1}^{n}f_{i}\lambda_{i}$
$\leqslant$ $\epsilon\sum_{i\neq r}f_{i}\lambda_{i}^{2}+\epsilon^{-1}\sum f_{i}+\big{|}\sum_{i=1}^{n}f_{i}\lambda_{i}\big{|},\,\forall\epsilon>0.$
@@
(8.21) $\mathcal{L}\widetilde{\Psi}\geqslant$ $A_{1}\sqrt{K}F^{i\bar{j}}(\mathfrak{\underline{g}}_{i\bar{j}}-\mathfrak{g}_{i\bar{j}})+2A_{1}N\sqrt{K}F^{i\bar{j}}\sigma_{i}\sigma_{\bar{j}}$
@@
$-(C_{1}^{\prime}+A_{3}C_{\Phi})\left(1+\big{|}\sum f_{i}\lambda_{i}\big{|}\right)-C_{1}^{\prime}\sum_{i=1}^{n}f_{i}.$
@@
Obviously $\widetilde{\Psi}(0)=0$. Using (8.16), (8.18) and $\mathcal{T}(u-\varphi)\big{|}_{\partial M}=0$ we can infer that $\widetilde{\Psi}\big{|}_{\partial{\Omega_{\delta}}}\leqslant 0$.
Next we will prove (8.22). The discussion can be divided into two cases according to Lemma 3.3.
@@
$F^{i\bar{j}}(\underline{\mathfrak{g}}_{i\bar{j}}-\mathfrak{g}_{i\bar{j}})\geqslant\varepsilon+\varepsilon F^{i\bar{j}}g_{i\bar{j}}+\varepsilon\big{|}F^{i\bar{j}}\mathfrak{g}_{i\bar{j}}\big{|}.$
@@
$F^{i\bar{j}}\sigma_{i}\sigma_{\bar{j}}\geqslant\frac{\varepsilon}{16}\left(1+\sum_{i=1}^{n}f_{i}+\big{|}\sum_{i=1}^{n}f_{i}\lambda_{i}\big{|}\right)\mbox{ on }\Omega_{\delta}$
@@
$\mathcal{L}\widetilde{\Psi}\geqslant$ $\frac{A_{1}N\varepsilon}{8}\sqrt{K}\left(1+\sum_{i=1}^{n}f_{i}+\big{|}\sum_{i=1}^{n}f_{i}\lambda_{i}\big{|}\right)-(C_{1}^{\prime}+A_{3}C_{\Phi})(1+\big{|}\sum_{i=1}^{n}f_{i}\lambda_{i}\big{|})$
$-C_{1}^{\prime}\sum_{i=1}^{n}f_{i}-\left(A_{2}+A_{3}C_{\Phi}+\frac{A_{1}\varepsilon}{32}+4(C_{1}^{\prime}+A_{3}C_{\Phi})^{2}\right)\sqrt{K}\sum_{i=1}^{n}f_{i}>0$